A piece of Fe metal with a mass of 0.350 kg at a temperature of 300 °C is added to 1.00 kg of H2O(1) at a temperature of 20 °C. What is the final equilibrium temperature of the H:0(0) and the Fe? The specific heat of the Fe metal is 0.502 J/g.°C and the specific heat of Hh0 is 4.18 Jlg. C.
Answer #1

Formula is as bellow

Q = mCp(T1 - Tf) where m = mass, Cp = Specific heat, T1 = Initial temprature while Tf = final temprature, Q = heat transfer. Now the heat lost by iron = heat gain by water

So the final eqution will be

0.35*0.502*(300-Tf) = 1*4.18*(Tf-20)

Or, 0.175*(300-Tf) = 4.18*(Tf-20)

Or, 300 + 477.71 = 24.885Tf

Or,  Tf = 31.25o C

Hence the final equibrium temprature of both of them will be 31.25o C